# Analytical Characterization of the Operational Diversity Order in Fading Channels

### v1

@@ IV-C When conventional diversity order is not applicable
$\delta_{\rm Ray}=\frac{W_{\rm th}}{\Omega_{\rm 0}}\frac{e^{-\tfrac{W_{\rm th}}{\Omega_{\rm 0}}}}{1-e^{-\tfrac{W_{\rm th}}{\Omega_{\rm 0}}}}.$ (17)
While for the latter the ODO quickly converges to unity, the former has a much slower growth and never manages to achieve the same diversity order as the single Rayleigh link. This is also represented in Fig. 5, where different linear approximations to the log-CDF are represented, using the slope values determined by the ODO. We see that assuming that the cascaded Rayleigh approximately decays with unitary slope may be exceedingly optimistic, and the true decay captured by the ODO must be considered for a proper system design around a given operation point.
Figure 4: ODO for the cascaded Rayleigh fading channel, as a function of the average SNR $\Omega$. Solid lines are obtained from (16), markers correspond to MC simulations. The case of the single Rayleigh fading channel is included as a reference.
Figure 5: ODO-based linear approximations to the OP for the cascaded Rayleigh fading channel at different operational points. Parameter values are $R=1.7$ bps/Hz. Solid lines correspond to the theoretical expressions for the OP, markers correspond to MC simulations.
@@ V Conclusions
We presented and formally characterized the operational diversity order, as an extension to the classical definition of diversity order but valid for any range of operational SNR. The ODO provides precise information about the decay slope of probability measures for communication systems operating over wireless fading channels, giving insightful information about how power excursions around the system’s SNR can impact its performance. Thanks to the ODO, the increased diversity order offered by LoS propagation is quantified, and the true diversity order in keyhole channels is analytically computed. The simple and compact definition of the ODO in terms of the aggregate fading PDF and CDF opens up the possibility to directly evaluate this insightful performance metric for any wireless system.
Proof of Lemma 1
@@
$=-\Omega\left.\frac{\partial\log P_{\rm op}\left(\Omega;\Gamma\right)}{\partial\Omega}\right|_{\Omega=\Omega_{\rm 0}},$ (21)
$=-\frac{\Omega}{\text{F}_{g}\left(W_{\rm th}/\Omega\right)}\left.\frac{\partial P_{\rm op}\left(\Omega;\Gamma\right)}{\partial\Omega}\right|_{\Omega=\Omega_{\rm 0}}.$ (22)
From (2), and applying the chain rule in the integral definition of the OP, we have

### v3

@@ IV-D When conventional DO is not applicable
$\delta_{\rm Ray}=\tfrac{W_{\rm th}}{\Omega_{\rm 0}}\tfrac{e^{-\tfrac{W_{\rm th}}{\Omega_{\rm 0}}}}{1-e^{-\tfrac{W_{\rm th}}{\Omega_{\rm 0}}}}.$ (16)
Figure 6: ODO-based linear approximations to the OP for the single and cascaded Rayleigh fading channels at different operational points (indicated by the colored $\times$ markers). Parameter values are $R=1.7$ bps/Hz. Solid lines correspond to the theoretical expressions for the OP.
While for the latter the ODO quickly converges to unity, the former has a much slower growth and never manages to achieve the same DO as the single Rayleigh link. This is also represented in Fig. 6, where different linear approximations to the log-CDF are represented, using the slope values determined by the ODO. We see that assuming that the cascaded Rayleigh approximately decays with unitary slope may be overly optimistic (e.g., slope is below 0.5/0.75 for $\Omega_{\rm 0}$ below 8/20 dB, respectively). Hence, the true decay captured by the ODO must be considered for a proper system design around a given operation point.
@@ V Conclusions
We presented and formally characterized the operational diversity order, as an extension to the classical definition of diversity order but valid in the vicinity of any operational SNR. Thanks to the ODO, the increased diversity order offered by LoS propagation is quantified, and the true diversity order in keyhole channels is analytically computed. The simple and compact definition of the ODO opens up the possibility to directly evaluate this insightful performance metric for any wireless system. Establishing the necessary and sufficient conditions for the ODO to exceed the conventional DO remains an open problem that requires further attention.
Appendix A Proof of Lemma 1
@@
=${\color[rgb]{0,0,0}\stackrel{{\scriptstyle(a)}}{{=}}}-\Omega\left.\frac{\partial\log P_{\rm op}\left(\Omega;\Gamma\right)}{\partial\Omega}\right|_{\Omega=\Omega_{\rm 0}},$ (20)
=${\color[rgb]{0,0,0}\stackrel{{\scriptstyle(b)}}{{=}}}-\frac{\Omega}{\text{F}_{g}\left(W_{\rm th}/\Omega\right)}\left.\frac{\partial P_{\rm op}\left(\Omega;\Gamma\right)}{\partial\Omega}\right|_{\Omega=\Omega_{\rm 0}}.$ (21)
where $(a)$ and $(b)$ follow after applying the chain rule.
From (2), and applying again the chain rule in the integral definition of the OP, we have
@@
Appendix B ODO under SC reception
Let $W_{\rm SC}=g_{\rm SC}\Omega$ be the instantaneous received power after SC in an $N$-branch receiver with i.i.d. branches [16], with $\Omega$ representing the average received power per branch. Noting that $F_{g_{\rm SC}}(x)=F_{g}(x)^{N}$ and $f_{g_{\rm SC}}(x)=Nf_{g}(x)F_{g}(x)^{N-1}$, the following identity holds:
${\color[rgb]{0,0,0}\delta_{\rm SC}=\frac{W_{\rm th}}{\Omega_{\rm 0}}\frac{f_{g_{\rm SC}}\left(W_{\rm th}/\Omega_{\rm 0}\right)}{{\rm F}_{g_{\rm SC}}\left(W_{\rm th}/\Omega_{\rm 0}\right)}=N\delta.}$ (23)